OpenStudy (anonymous):

Solve for x. 4x(x-9)-5x(x-8)=-5

sam (.sam.):

4x(x-9)-5x(x-8)=-5 Multiply 4x by each term inside the parentheses. 4x^(2)-36x-5x(x-8)=-5 Multiply -5x by each term inside the parentheses. 4x^(2)-36x-5x^(2)+40x=-5 Since 4x^(2) and -5x^(2) are like terms, add -5x^(2) to 4x^(2) to get -x^(2). -x^(2)-36x+40x=-5 Since -36x and 40x are like terms, subtract 40x from -36x to get 4x. -x^(2)+4x=-5 To set the left-hand side of the equation equal to 0, move all the expressions to the left-hand side. -x^(2)+4x+5=0 Multiply each term in the equation by -1. x^(2)-4x-5=0 In this problem 1*-5=-5 and 1-5=-4, so insert 1 as the right hand term of one factor and -5 as the right-hand term of the other factor. (x+1)(x-5)=0 The complete solution is the set of the individual solutions. x=-1,5

OpenStudy (anonymous):

4x(x-9)-5x(x-8)=-5 4x² - 36x - 5x² + 40x = -5 -x² + 4x + 5 = 0 x² - 4x - 5 = 0 4 +- sqrt 36 / 2 4 +- 6 / 2 10/2 and -2/2 x = 5 and x = -1